OpenStudy (anonymous):

Solve for y: x=(2y-3)/(y+1)

OpenStudy (anonymous):

xy+x=2y-3 x+3=2y-xy x+3=y(2-x) y=(x+3)/(2-x)

OpenStudy (anonymous):

x neq 0